OpenStudy (anonymous):

A cruise ship can cover 18 nautical miles in 324 minutes. How many nautical miles will it travel in 180 minutes? A. 12 B. 14 C. 8 D. 10

OpenStudy (anonymous):

@AihberKhan

OpenStudy (wolf1728):

Set up a proportion 18/324 = x/180 Can you solve for x?

OpenStudy (jdoe0001):

\(\bf \begin{array}{ccllll} miles&minutes \\\hline\\ 18&324\\ x&180 \end{array}\implies \cfrac{18}{x}=\cfrac{324}{180}\) solve for "x"

OpenStudy (anonymous):

thanks

OpenStudy (wolf1728):

u r welcome